OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5.

OpenStudy (texaschic101):

first of all, parallel lines will have the same slope. So we need to find the slope of 7x + y = -5. Do you know how to find the slope of this line ?

OpenStudy (anonymous):

m=-7

OpenStudy (texaschic101):

right.....slope is -7...parallel lines will have same slope now we will use y = mx + b slope (m) = -7 (1,-3)....x = 1 and y = -3 we have both the slope and our points, so we basically need to find b, the y intercept. y = mx + b -3 = -7(1) + b -3 = -7 + b -3 + 7 = b 4 = b therefore, our parallel line is : y = -7x + 4 questions ?

OpenStudy (anonymous):

thanks (:

OpenStudy (texaschic101):

anytime :)